# Quantum Zeno Effect in Noisy Integrable Quantum Circuits for Impurity Models

### v1

@@ Abstract
We theoretically study the open quantum system dynamics (in the Trotterized limit) of integrable quantum circuits in the presence of onsite dephasing noise with a spin-1/2 “impurity” interacting at the edge. Using a combination of Bethe Ansatz (BA) and exact diagonalization (ED), we study the dynamics of both the bulk and the impurity for the XXX (Heisenberg) and the XX qubit chains in the presence and absence of bulk noise. In the absence of noise, we show that the impurity exhibits two distinct phases, the bound mode phase where the impurity keeps oscillating in time, and the Kondo phase where it decays with Kondo time $t_{K}$. Turning on the bulk dephasing noise, we find for the two models that in the long time limit in both regimes the quantum Zeno effect takes place where the dynamics of the impurity magnetization slows down as the noise strength $\gamma$ increases. The impurity magnetization in the bound mode regime shows the opposite effect, decaying faster as the noise strength increases for short times ($t\ll 1/\gamma$). We show that the bulk KPZ dynamics of the XXX model is converted to diffusive dynamics as in the XX case studied before by V. Alba, driving both systems to the Zeno effect slowing down the impurity dynamics in the long time limit.
Introduction: The rise of noisy intermediate scale quantum (NISQ) [1, 2, 3] devices in a wide range of physical platforms including superconducting qubits [4, 5, 6], trapped ions [7, 8, 9, 10], neutral atoms [11, 12, 13], and circuit quantum electrodynamics [14, 15, 16, 17, 18], has made understanding the effects of noise on these devices and on quantum many-body systems (QMBS) in general a question of paramount importance [19, 20, 21, 22]. Simulating large generic QMBS on a classical computer is extremely difficult  [23, 24, 25] as the size of the Hilbert space grows exponentially with the size of the system. To circumvent this challenge, quantum computers can be utilized to study QMBS [24, 25] (as they are built to hold such an exponentially large Hilbert space). This can be achieved by Trotterizing the continuous time evolution of a QMBS as a quantum circuit [26, 27]. While, generically solving a quantum circuit remains an equally challenging problem, integrable quantum circuits [28, 29, 30, 31, 32, 33], in particular, based on the $R$-matrices that solve the Yang-Baxter equation  [34, 35], can simulate the unitary evolution of initial states under integrable Hamiltonians [36, 37, 38, 39]. Integrability allows more control and insights into the behavior of these complex systems garnering significant attention recently  [40, 41, 42].
All quantum devices currently accessible are under a significant influence of noise [43, 44, 45, 46, 47, 48, 49, 50, 51, 52]. Noise affects the dynamics and transport properties of quantum circuits resulting in dramatic changes [19, 20, 21]. Interesting effects arise such as dynamical phase transitions [53, 54], the quantum Zeno effect (where increasing the strength of noise slows down quantum dynamics [55, 56]) and the quantum Mpemba effect (where non-equilibrium quantum systems relax faster when they are further from equilibrium [57, 58, 59, 60]). Thus, studying the effect of noise on quantum circuits is of crucial importance to understand the experiments performed on NISQ devices [61]. Here, we study a circuit that can be solved exactly in the presence of dephasing noise.
@@
Models: To begin with, we construct an integrable quantum circuit that corresponds to a unitary operator
@@
where $L$ is the number of qubits, $\check{R}_{j,j+1}$ given explicitly in [62] is the braid form of the trigonometric $R$-matrix of the six-vertex model [63, 36] acting on the $j$ and $j+1$ qubits, $K_{01}^{\eta}(\lambda,d)=\check{R}_{01}^{\eta}(-d)\check{R}^{\eta}_{01}(\lambda+d)$ is the boundary $K$-matrix that solves the reflection algebra [64, 65, 66], and $d$, $\lambda$ and $\eta$ are free parameters. In the Trotterziation limit [67, 68], choosing $\lambda=-i\delta t\sinh\eta$, this quantum circuit simulates the unitary evolution $\exp(-iHt)=\lim_{\delta t\to 0}\mathbb{U}(\lambda,d,\eta)^{\frac{t}{\delta t}}$  for an integrable Hamiltonian $H$ [25, 27, 69], where
$\begin{split}H=&\sum_{j=1}^{L-1}\frac{1}{2}(\sigma_{j}^{x}\sigma_{j+1}^{x}+\sigma_{j}^{y}\sigma_{j+1}^{y}+\Delta\sigma_{j}^{z}\sigma_{j+1}^{z})\\ &+\frac{J}{2}\left(\sigma_{0}^{x}\sigma_{1}^{x}+\sigma_{0}^{y}\sigma_{1}^{y}+\Delta^{\prime}\sigma_{0}^{z}\sigma_{1}^{z}\right)\\ \end{split}$ (2)
with $\Delta\mathord{=}\cosh(\eta)$, $J\mathord{=}\frac{\sinh^{2}(\eta)\cosh(d)}{\sinh^{2}(\eta)-\sinh^{2}(d)}$ and $\Delta^{\prime}\mathord{=}\frac{\cosh(\eta)}{\cosh(d)}$. This Hamiltonian describes the $XXZ$ spin chain with a boundary impurity. At the isotropic XXX ($\Delta\mathord{=}\Delta^{\prime}\mathord{=}1)$ and the XX ($\Delta\mathord{=}\Delta^{\prime}\mathord{=}0$) points, the model is integrable for arbitrary boundary coupling $J$. For generic $\Delta$, the boundary coupling ($J$) and the boundary anisotropy parameter ($\Delta^{\prime}$) are related and expressed in terms of a single free parameter $d$ to maintain the integrability [70, 71, 72, 73, 74, 75, 76]. In the ground state, the impurity is screened due to the Kondo effect when $\bar{d}\mathord{=}d/\eta$ is purely imaginary or $0<\bar{d}<\frac{1}{2}$ and it is screened by a bound mode exponenitially localized at the boundary when $1>\bar{d}>\frac{1}{2}$ [70].
@@
where the jump operators are $L_{j}\mathord{=}\sigma^{z}_{j}$, $\gamma$ is the dephasing rate, $H$ is the Hamiltonian in Eq. (2), $\{A,B\}$ and $[A,B]$ denote the anticommutator and the commutator, respectively. In the following, we will evolve the system from two initial states, Néel state and domain wall state expressed as: $\rho_{N/D}=(\prod_{z\in\mathcal{S}_{N/D}}\sigma^{+}_{z})\ket{0}\bra{0}(\prod_{z\in\mathcal{S}_{N/D}}\sigma^{-}_{z})$ where $\mathcal{S}_{N/D}$ denote the set of sites with $z\text{ }\mathord{=}\text{ }0,2,4,...,L$ for the Néel (N) state $\rho_{N}$, and $z\text{ }\mathord{=}\text{ }0,1,2,...,L/2$ for the domain wall (D) state $\rho_{D}$. Here, |$\ket{0}$ refers to the state with all spins down (qubits 00) state. In both cases, we compute the time-dependent expectation value of magnetization at site $j$, $S^{z}_{j}(t)\mathord{=}\mathrm{Tr}\{\sigma^{z}_{j}\rho(t)\}$.
@@
Figure 1: (a) The parameter space $(J,\gamma)$ is divided into two distinct regimes: the bound-mode regime, and the Kondo regime. The real part of the double-bound mode energy is computed for $L=30$ is given in the heatmap. For large $\gamma\mathord{\gg}1$, the boundary between two regimes is almost $J\mathord{=}\gamma/2$ as shown in the dashed line in (a). The Liouvillian spectrum and amplitude $\alpha_{E}=G_{E}(0,0)(\bar{G}_{E}|G_{in})$ for Néel state in the two-particle sector with dephasing rate $\gamma=0.25$ and impurity coupling $J=0.5$ (b) in the Kondo regime, $\gamma=2$, $J=0.5$ (c) in the Kondo regime, and $\gamma=2$, $J=3$ (d) in the bound mode regime due to the existence of the double-bound mode. The data points labeled in the legend with different shapes and numbers represent four different kinds (respectively correspond to the four kinds of solutions solved analytically in the main text). The scattering color plots are numerical results solved with system size $L=40$ where the heatmap shows absolute values of amplitude $\alpha_{E}^{N}$ for each eigenmode. The modes with larger $\alpha_{E}^{N}$ dominate the impurity dynamics: the double-bound mode in the bound mode regime (d), and the modes within the diffusive band in the Kondo regime (b, c). With a small noise $\gamma<1$, the diffusive band is mixed with the other bands (b), and when $\gamma>1$ the diffusive band is separated from the other bands (c,d).
Figure 2: The dynamics of the XX and XXX impurity models evolving from a Néel initial state under a Lindbladian evolution with and without dephasing noise. Different colors represent different dephasing rates $\gamma$, solid lines for $J=0.5$ in the Kondo regime, and dashed lines for $J=3$ in the bound-mode regime. The upper panel shows the data for the XXX chain for $L=10$ and the bottom panel for the XX chain for $L=40$. For both XX and XXX chains, in the bound mode phase, the signal decays faster as the dephasing rate increases. However, in the Kondo phase, the impurity decay is slower as the dephasing rate increases. The insets in each panel show the data on a log-linear scale displaying the same power law decay independent of $J$ in the long-time limit.
We now proceed to solve for the spectrum $E$ of the Liouvillian $\hat{h}$ defined in Eq. (4). In the absence of the impurity, this Liouvillian can be mapped to an integrable non-Hermitian Hubbard model after a unitary transformation [81, 82, 83, 84, 85], here we generalize it to include the impurity as discussed in [62]. The Liouvillian in Eq. (4) can thus be solved via the Bethe Ansatz (BA) [86, 87, 88]. It has eigenvalue $E(k,q)=2\cos k+2\cos q-4i\gamma$ with the eigenmodes wavefunction taking the BA form
@@
If $(k,q)$ is a solution of the BAE (8), then (±$(\pm k+2\pi,\pm q+2\pi)$ is also a solution. This allows us to restrict the Brillouin zone to $\Re k\in(0,\pi)$ with $\Re k$ ($\Im k$) as the real (imaginary) part of $k$. Furthermore, its complex conjugate $(k^{*},q^{*})$ also being a solution, indicates that if $E$ is a Liouvillian eigenvalue, then $-E^{*}$ is also an eigenvalue [82]. The quasimomenta $(k,q)$ satisfying the BAEs fall into one of the following four cases in the limit $L\to\infty$. Here, we briefly summarize the resuls and details can be found in [62].
@@
1.Free band: When both S-matrices are regular $S,K\neq 0,\infty$, the quasimomenta $k$ and $q$ are complex, with imaginary parts vanishing as $\sim\frac{1}{L}$, i.e $k=\frac{\pi n_{1}}{L+1}+\mathcal{O}(L^{-1})$ and $q=\frac{\pi n_{2}}{L+1}+\mathcal{O}(L^{-1})$, for $n_{1},n_{2}=1,\ldots,L$. These solutions exist for all parameters.
@@
2. Bound mode band: When $K\mathord{=}\infty$, quasimomenta are $k_{\pm}=i\log\left(\frac{\pm\sqrt{J^{2}-1-\gamma^{2}}-i\gamma}{J^{2}-1}\right)$ and $q=\frac{\pi n}{L+1}+\mathcal{O}(L^{-1})$ for $n=1,\ldots,L$. If $\Im k_{\pm}<0$, it is a bound mode; otherwise, it is not a valid solution. When both $\Im k_{\pm}<0$, another mode with $E=2\cos k_{+}+2\cos k_{-}-4i\gamma$ exists.
@@
3. Diffusive band: When $S\mathord{=}\infty$, the eigenspectra are $E_{n}=4i\sqrt{\gamma^{2}-\sin\left(\frac{\pi n}{2L}\right)^{2}}-4i\gamma+\mathcal{O}(L^{-3})$ with $n=0,1,2,\ldots,L$. Gapless excitations with $E_{n}=-i\frac{\pi^{2}n^{2}}{2\gamma L^{2}}+\mathcal{O}(L^{-3})$ [81, 82], dominate the late time dynamics. These solutions with $\Re E_{n}=0$ exist for all parameters.
@@
4. Double-bound mode: There are two solutions where the particles are boundary-localized. When $\gamma\mathord{<}1$, the eigenvalues can be solved perturbtively $E_{\pm}=4\cos k_{\pm}+4i\gamma\left[\sum_{x=1}^{L}\mathcal{N}^{2}e^{i4k_{\pm}x}+\frac{\mathcal{N}}{J^{2}}-1\right]+\mathcal{O}(\gamma^{2})$, where $k_{\pm}=\frac{i}{2}\log(J^{2}-1)+\pi/2\pm\pi/2$ and $\mathcal{N}=\frac{1}{J^{2}}+\frac{e^{-2ik_{\pm}}}{1-e^{-2ik_{\pm}}}$. When $\gamma>1$, the eigenvalues solved with an effective two-site Liouvillian are $E_{\pm}=\pm\sqrt{4J^{2}-\gamma^{2}}-i\gamma$.
@@
Fig. 1 shows the spectrum of the Liouvillian in the Kondo and bound regimes. We present the cross-over diagram that delineates these two regimes in Fig. 1(a) through the value of the real part of the Double bound mode $\Re E_{DBM}$. A non-zero $\Re E_{DBM}$ signifies the bound mode regime, while a zero value indicates the Kondo regime. We further analyze the spectrum’s real and imaginary parts in the Kondo regime in Fig. 1(a) and (b) as well as the bound mode regime in Fig. 1 (c). We see the free band (spectrum 1; circles), the bound mode band (spectrum 2; crosses), and the diffusive band (spectrum 3; star) in the Kondo regime, where the majority of the weight in $\alpha_{E}^{N}$ lies in the diffusive band. In the bound mode regime, we also see the appearance of double bound modes (spectrum 4; diamonds) that have the largest weight in the coefficients $\alpha_{E}^{N}$ and hence they dominate the early-time impurity dynamics $S^{z}_{0}(t)$.
@@
In [71], we showed that at the XX point without noise, there is a boundary phase transition at $J=\sqrt{2}$. When $J<\sqrt{2}$, the impurity is screened by many-body Kondo effect whereas when $J>\sqrt{2}$, the impurity is screened by single particle bound mode localized at the edge. Here, we turn to study the dynamical characterization of the boundary phases. In the bound mode phase ($J>\sqrt{2}$), the time-dependent magnetization shows oscillatory behavior $S^{z}_{0}(t)\sim\cos(\frac{J^{2}}{\sqrt{J^{2}-1}}t)$ when $t\to\infty$. However in the Kondo phase ($J<\sqrt{2}$), the impurity magnetization leaks to the bulk with the time scale proportional to the Kondo time $t_{K}=\sqrt{|J^{2}-1|}/J^{2}$. Detailed calculation showing this sharp phase transition in dynamical quantities is in [62].
@@
When noise is introduced ($\gamma>0$), there is no longer a sharp phase transition. Instead, the impurity magnetization $S^{z}_{0}(t)$ decays similarly in both regimes at late times, exhibiting the Zeno effect, where the dynamics slows down as noise strength increases. This behavior is driven by the diffusive band as it has largest overlap $\alpha_{E}^{N}$. The modes in the diffusive band have $|\Im E(\gamma)|\propto\gamma^{-1}$, so the dynamics slows down as $\gamma$ increases and exhibits the quantum Zeno effect. In contrast, the early-time dynamics differ between the two regimes. In the bound mode regime, the impurity dynamics is dominated by the double-bound mode which has the largest overlap $\alpha_{E}^{N}$ and $|\Im E(\gamma)|\propto\gamma$. Thus, the impurity decays faster as $\gamma$ increases. However, in the Kondo regime, the early-time dynamics are also influenced by the diffusive band—similar to the long-time limit where the decay rate decreases with increasing noise, resulting the quantum Zeno effect. The model parameters $J$ and $\gamma$ corresponding to these two regimes are shown in Fig. 1(a).
@@
Numerical results for $S^{z}_{0}(t)$ starting from the Néel state at the XX point are shown in the bottom panel of Fig. 2 with two situations when the system is in the Kondo regime $J=0.5$ and in the bound-mode regime $J=3$. As seen from the data, for the early time in the bound-mode regime, the noise adds a decaying envelope that is accelerated as the dephasing rate is increased. This is an expected result because the boundary mode is essentially a two-qubit state, and qubits subjected to dephasing noise decay faster as the decay rate increases. In the Kondo regime or after the double bound mode decays i.e. $t\sim\gamma^{-1}$, the noise slows the leakage of magnetization and leads to a quantum Zeno effect as the dephasing rate increases. In the long time limit, independent of the boundary coupling $J$, the impurity magnetization falls off as $S^{z}_{0}(t)\sim e^{-t/\tau}$ with $\tau\propto\frac{1}{\gamma L^{2}}$ in both regimes. This behavior of the impurity in the Kondo regime is consistent with the heating of the bulk induced by Hermitian Lindblad operators [89, 90, 91], driving the impurity to (asymptotically free) weak coupling behavior [92, 93, 94, 95].
@@
Figure 3: Data collapse for $S^{z}(x,t)=S^{z}(\xi)$ with $\xi=xt^{-\frac{1}{2}}$ in the XX case (filled markers, with system size $L=40$) and XXX case (hollow markers, with system size $L=10$) starting from a domain wall initial state (as shown in the color plot in the right inset for XX case with $\gamma=2$) under various dephasing rates $\gamma$. Different marker shapes correspond to the spin configuration at different time slices, and different marker color refers to different dephasing rates. The left inset shows the dependence of the diffusion constant $D$ on $\gamma$ (filled markers for XX case and hollow for XXX case) and the dashed lines are the analytical result for the noisy XX chain $D=\frac{1}{2\gamma}$.

### v2

@@ Abstract
We theoretically study the open quantum system dynamics (in the Trotterized limit) of integrable quantum circuits in the presence of onsite dephasing noise with a spin-1/2 “impurity” interacting at the edge. Using a combination of Bethe Ansatz (BA) and exact diagonalization (ED), we study the dynamics of both the bulk and the impurity for the XXX (Heisenberg) and the XX qubit chains in the presence and absence of bulk noise. In the absence of noise, we show that the impurity exhibits two distinct phases: the bound mode phase, where the impurity keeps oscillating in time, and the Kondo phase, where it decays with Kondo time $t_{K}$. Turning on the bulk dephasing noise, we find for the two models that in the long-time limit in both regimes, the quantum Zeno effect takes place where the dynamics of the impurity magnetization slows down as the noise strength $\gamma$ increases. The impurity magnetization in the bound-mode regime shows the opposite effect, decaying faster as the noise strength increases for short times ($t\ll 1/\gamma$). We show that the Zeno effect slowing down the impurity dynamics in the long-time limit is driven by the change in bulk dynamics from ballistic (KPZ) universality class in the XX (XXX) chain to diffusive dynamics in the presence of noise.
Introduction: The interplay between noise and quantum coherence is central to understanding modern quantum devices, particularly in noisy intermediate-scale quantum (NISQ) platforms such as superconducting qubits [1, 2, 3], trapped ions [4, 5, 6, 7], neutral atoms [8, 9, 10], and circuit quantum electrodynamics [11, 12, 13, 14, 15]. In these open systems [16, 17, 18, 19], noise—ranging from dephasing and relaxation to control errors—fundamentally limits coherence times and gate fidelities [20, 21], making it imperative to study its effects on quantum many-body systems (QMBS). Classically simulating large QMBS is infeasible due to exponential Hilbert space growth [22, 23, 24], but quantum computers naturally accommodate such complexity [23, 24]. A promising approach involves Trotterizing time evolution into quantum circuits [25, 26], with integrable circuits [27, 28, 29, 30, 31, 32] offering unique advantages. Built from Yang-Baxter-solvable $R$-matrices [33, 34], these circuits simulate integrable Hamiltonians [35, 36, 37, 38] with exact control over dynamics, enabling insights into complex many-body behavior [39, 40, 41, 42].
However, real-world quantum devices are inherently noisy [43, 44, 45, 46, 47, 48, 49, 50, 51, 52], which disrupts transport and induces phenomena like dynamical phase transitions [53, 54], the quantum Zeno effect (where increasing the strength of noise slows down quantum dynamic) [55, 56, 57], and the quantum Mpemba effect (where non-equilibrium quantum systems relax faster when they are further from equilibrium) [58, 59, 60, 61]. To disentangle these effects, we study an exactly solvable integrable circuit with dephasing noise, bridging the gap between ideal unitary dynamics and realistic noisy environments. Our work reveals how noise inherently modifies coherent evolution, offering a controlled framework to benchmark NISQ experiments [62] by studying anisotropic Heisenberg spin chains which can be simulated on experimentally available platforms [63, 64].
@@
Models: To begin with, we construct an integrable quantum circuit as shown in Fig.1 that corresponds to a unitary operator
$\mathbb{U}^{\eta}(\lambda,d)$
Figure 1: Illustrations of an integrable quantum circuit with spin-1/2 qubits in an initial classical Néel state (empty circles are qubits in the up states and black filled circles are in the down qubit state). Each rectangle represents a local unitary matrix; blue ones are the braid form of the trigonometric six-vertex R-matrix $\check{R}(\lambda)$ and red ones are the boundary reflection matrix $K(\lambda,d)=\check{R}(-d)\check{R}(\lambda+d)$. The dashed rectangle shows one layer of the time evolution operator $\mathbb{U}^{\eta}(\lambda,d)$.
@@
where $L$ is the number of qubits, $\check{R}_{j,j+1}$ given explicitly in Appendix A is the braid form of the trigonometric $R$-matrix of the six-vertex model [65, 35] acting on the $j$ and $j+1$ qubits, $K_{01}^{\eta}(\lambda,d)=\check{R}_{01}^{\eta}(-d)\check{R}^{\eta}_{01}(\lambda+d)$ is the boundary $K$-matrix that solves the reflection algebra [66, 67, 68], and $d$, $\lambda$ and $\eta$ are free parameters. In the Trotterziation limit [69, 70], choosing $\lambda=-i\delta t\sinh\eta$, this quantum circuit simulates the unitary evolution $\exp(-iHt)=\lim_{\delta t\to 0}\mathbb{U}(\lambda,d,\eta)^{\frac{t}{\delta t}}$  for an integrable Hamiltonian $H$ [24, 26, 71], where
[…]
with $\Delta\mathord{=}\cosh(\eta)$, $J\mathord{=}\frac{\sinh^{2}(\eta)\cosh(d)}{\sinh^{2}(\eta)-\sinh^{2}(d)}$ and $\Delta^{\prime}\mathord{=}\frac{\cosh(\eta)}{\cosh(d)}$. This Hamiltonian describes the $XXZ$ spin chain with a boundary impurity. At the isotropic XXX ($\Delta\mathord{=}\Delta^{\prime}\mathord{=}1)$ and the XX ($\Delta\mathord{=}\Delta^{\prime}\mathord{=}0$) points, the model is integrable for arbitrary boundary coupling $J$. For generic $\Delta$, the boundary coupling ($J$) and the boundary anisotropy parameter ($\Delta^{\prime}$) are related and expressed in terms of a single free parameter $d$ to maintain the integrability [72, 73, 74, 75, 76, 77, 78]. In the ground state, the impurity is screened due to the Kondo effect when $\bar{d}\mathord{=}d/\eta$ is purely imaginary or $0<\bar{d}<\frac{1}{2}$ and it is screened by a bound mode exponentially localized at the boundary when $1>\bar{d}>\frac{1}{2}$ [72].
@@
XX chain: Let us start from the XX point as it allows for a simpler treatment with similar phenomenology [73]. It is convenient to work in the fermionic language via a Jordan-Wigner transformation $c^{\dagger}_{j}\mathord{=}\sigma^{+}_{j}\prod_{l<j}\sigma^{z}_{l}$. The magnetization is then related to the two-point correlator $G(x,y;t)=\mathrm{Tr}[c^{\dagger}_{x}c_{y}\rho(t)]$ as $S^{z}_{j}(t)=2G(j,j;t)-1$. A closed set of equations gives the dynamics of the two-point correlators because the Hamiltonian and the jump operators are both quadratic in fermions [83, 84]. Therefore, the dynamics of the two-point correlation function in the noisy XX model can be mapped to a two-particle quantum mechanical problem as shown in Appendix B. The correlator can be vectorized as a state (see Appendix B) $|G(t))\in\mathbb{C}^{L^{2}}$ where $|G(t))=\sum_{x,y}G(x,y;t)|x,y)$ and $|x,y)$ forms an orthonormal basis in $\mathbb{C}^{L^{2}}$. Its evolution from an initial state $|G_{in})=\sum_{x,y}\mathrm{Tr}[G(x,y{;0})\rho_{N/D}]|x,y)$ reads $|G(t))=e^{-i\hat{h}t}|G_{in})$ with Liouvillian $\hat{h}=\sum_{x,y,x^{\prime},y^{\prime}}h^{x,y}_{x^{\prime},y^{\prime}}|x,y)(x^{\prime},y^{\prime}|$ and
@@
Figure 2: (a) The parameter space $(J,\gamma)$ is divided into two distinct regimes: the bound-mode regime and the Kondo regime. The real part of the double-bound-mode energy is computed for $L=30$ and is given in the heatmap. For large $\gamma\mathord{\gg}1$, the boundary between two regimes is almost $J\mathord{=}\gamma/2$ as shown in the dashed line in (a). The Liouvillian spectrum and amplitude $\alpha_{E}=G_{E}(0,0)(\bar{G}_{E}|G_{in})$ for Néel state in the two-particle sector with dephasing rate $\gamma=0.25$ and impurity coupling $J=0.5$ (b) in the Kondo regime, $\gamma=2$, $J=0.5$ (c) in the Kondo regime, and $\gamma=2$, $J=3$ (d) in the bound-mode regime due to the existence of the double-bound mode. The data points labeled in the legend with different shapes and numbers represent four different kinds (respectively correspond to the four kinds of solutions solved analytically in the main text). The scattering color plots are numerical results solved with system size $L=40$ where the heatmap shows absolute values of amplitude $\alpha_{E}^{N}$ for each eigenmode. The modes with larger $\alpha_{E}^{N}$ dominate the impurity dynamics: the double-bound mode in the bound-mode regime (d) and the modes within the diffusive band in the Kondo regime (b, c). With a small noise $\gamma<1$, the diffusive band is mixed with the other bands (b), and when $\gamma>1$, the diffusive band is separated from the other bands (c,d).
Figure 3: The dynamics of the XX and XXX impurity models evolving from a Néel initial state under a Lindbladian evolution with and without dephasing noise. Different colors represent different dephasing rates $\gamma$, solid lines for $J=0.5$ in the Kondo regime, and dashed lines for $J=3$ in the bound-mode regime. The upper panel shows the data for the XXX chain for $L=10$ and the bottom panel for the XX chain for $L=40$. For both XX and XXX chains, in the bound mode phase, the signal decays faster as the dephasing rate increases. However, in the Kondo phase, the impurity decay is slower as the dephasing rate increases. The insets in each panel show the data on a log-linear scale, displaying the same power law decay independent of $J$ in the long-time limit.
We now proceed to solve for the spectrum $E$ of the Liouvillian $\hat{h}$ defined in Eq. (4). In the absence of the impurity, this Liouvillian can be mapped to an integrable non-Hermitian Hubbard model after a unitary transformation [84, 85, 86, 87], here we generalize it to include the impurity as discussed in Appendix A. The Liouvillian in Eq. (4) can thus be solved via the Bethe Ansatz (BA) [88, 89, 90]. It has eigenvalue $E(k,q)=2\cos k+2\cos q-4i\gamma$ with the eigenmodes wavefunction taking the BA form
@@
If $(k,q)$ is a solution of the BAE (8), then $(\pm k+2\pi,\pm q+2\pi)$ is also a solution. This allows us to restrict the Brillouin zone to $\Re k\in(0,\pi)$ with $\Re k$ ($\Im k$) as the real (imaginary) part of $k$. Furthermore, its complex conjugate $(k^{*},q^{*})$ also being a solution, indicates that if $E$ is a Liouvillian eigenvalue, then $E^{*}$ is also an eigenvalue [85]. The quasimomenta $(k,q)$ satisfying the BAEs fall into one of the following four cases in the limit $L\to\infty$. Here, we briefly summarize the results, and details can be found in Appendix D.
@@
1.Free band: When both S-matrices are regular $S,K\neq 0,\infty$, the quasimomenta $k$ and $q$ are complex, with imaginary parts vanishing as $\sim\frac{1}{L}$, i.e $k=\frac{\pi n_{1}}{L+1}+\mathcal{O}(L^{-1})$ and $q=\frac{\pi n_{2}}{L+1}+\mathcal{O}(L^{-1})$, for $n_{1},n_{2}=1,\ldots,L$. These solutions exist for all parameters. These solutions dominate when $t\ll(4\gamma)^{-1}$ and generate noiseless dynamics.
@@
2. Bound mode band: When $K\mathord{=}\infty$, quasimomenta are $k_{\pm}=i\log\left(\frac{\pm\sqrt{J^{2}-1-\gamma^{2}}-i\gamma}{J^{2}-1}\right)$ and $q=\frac{\pi n}{L+1}+\mathcal{O}(L^{-1})$ for $n=1,\ldots,L$. If $\Im k_{\pm}<0$, it is a bound mode; otherwise, it is not a valid solution. When both $\Im k_{\pm}<0$, another mode with $E=2\cos k_{+}+2\cos k_{-}-4i\gamma$ exists. These solutions are similar as the free ones, but they generate bound-mode dynamics for early times.
@@
3. Diffusive band: When $S\mathord{=}\infty$, the eigenspectra are $E_{n}=4i\sqrt{\gamma^{2}-\sin\left(\frac{\pi n}{2L}\right)^{2}}-4i\gamma+\mathcal{O}(L^{-3})$ with $n=0,1,2,\ldots,L$. Gapless excitations with $E_{n}=-i\frac{\pi^{2}n^{2}}{2\gamma L^{2}}+\mathcal{O}(L^{-3})$ [92, 84, 85], dominate the late time dynamics. These solutions with $\Re E_{n}=0$ exist for all parameters, and they generate the diffusive dynamics due to the dephasing noise, known as Incoherentons[93].
@@
4. Double-bound mode: There are two solutions where the particles are boundary-localized. When $\gamma\mathord{<}1$, the eigenvalues can be solved perturbtively $E_{\pm}=4\cos k_{\pm}+4i\gamma\left[\sum_{x=1}^{L}\mathcal{N}^{2}e^{i4k_{\pm}x}+\frac{\mathcal{N}}{J^{2}}-1\right]+\mathcal{O}(\gamma^{2})$, where $k_{\pm}=\frac{i}{2}\log(J^{2}-1)+\pi/2\pm\pi/2$ and $\mathcal{N}=\frac{1}{J^{2}}+\frac{e^{-2ik_{\pm}}}{1-e^{-2ik_{\pm}}}$. When $\gamma>1$, the eigenvalues solved with an effective two-site Liouvillian are $E_{\pm}=\pm\sqrt{4J^{2}-\gamma^{2}}-i\gamma$. These solutions dominate the early-time oscillation in the magnetization in the bound-mode regime at short times.
@@
Fig. 2 shows the spectrum of the Liouvillian in the Kondo and bound regimes. We present the cross-over diagram that delineates these two regimes in Fig. 2(a) through the value of the real part of the Double bound mode $\Re E_{DBM}$. A non-zero $\Re E_{DBM}$ signifies the bound-mode regime, while a zero value indicates the Kondo regime. We further analyze the spectrum’s real and imaginary parts in the Kondo regime in Fig. 2(a) and (b), as well as the bound-mode regime in Fig. 2 (c). We see the free band (spectrum 1; circles), the bound mode band (spectrum 2; crosses), and the diffusive band (spectrum 3; star) in the Kondo regime, where the majority of the weight in $\alpha_{E}^{N}$ lies in the diffusive band. In the bound-mode regime, we also see the appearance of double bound modes (spectrum 4; diamonds) that have the largest weight in the coefficients $\alpha_{E}^{N}$ and hence they dominate the early-time impurity dynamics $S^{z}_{0}(t)$.
@@
In [73], we showed that at the XX point without noise, there is a boundary phase transition at $J=\sqrt{2}$. When $J<\sqrt{2}$, the impurity is screened by the many-body Kondo effect, whereas when $J>\sqrt{2}$, the impurity is screened by single particle bound mode localized at the edge. Here, we turn to study the dynamical characterization of the boundary phases. In the bound mode phase ($J>\sqrt{2}$), the time-dependent magnetization shows oscillatory behavior $S^{z}_{0}(t)\sim\cos(\frac{J^{2}}{\sqrt{J^{2}-1}}t)$ when $t\to\infty$. However, in the Kondo phase ($J<\sqrt{2}$), the impurity magnetization leaks to the bulk with the time scale proportional to the Kondo time $t_{K}=\sqrt{|J^{2}-1|}/J^{2}$. Detailed calculation showing this sharp phase transition in dynamical quantities is in Appendix C.
@@
When noise is introduced ($\gamma>0$), there is no longer a sharp phase transition. Instead, the impurity magnetization $S^{z}_{0}(t)$ decays similarly in both regimes at late times, exhibiting the Zeno effect, where the dynamics slows down as noise strength increases. This behavior is driven by the diffusive band as it has the largest overlap $\alpha_{E}^{N}$. The modes in the diffusive band have $|\Im E(\gamma)|\propto\gamma^{-1}$, so the dynamics slows down as $\gamma$ increases and exhibits the quantum Zeno effect. In contrast, the early-time dynamics differ between the two regimes. In the bound-mode regime, the impurity dynamics is dominated by the double-bound mode, which has the largest overlap $\alpha_{E}^{N}$ and $|\Im E(\gamma)|\propto\gamma$. Thus, the impurity decays faster as $\gamma$ increases. However, in the Kondo regime, the early-time dynamics are also influenced by the diffusive band—similar to the long-time limit where the decay rate decreases with increasing noise, resulting in the quantum Zeno effect. The model parameters $J$ and $\gamma$ corresponding to these two regimes are shown in Fig. 2(a).
@@
Numerical results for $S^{z}_{0}(t)$ starting from the Néel state at the XX point are shown in the bottom panel of Fig. 3 with two situations when the system is in the Kondo regime $J=0.5$ and in the bound-mode regime $J=3$. As seen from the data, for the early time in the bound-mode regime, the noise adds a decaying envelope that is accelerated as the dephasing rate is increased. This is an expected result because the boundary mode is essentially a two-qubit state, and qubits subjected to dephasing noise decay faster as the decay rate increases. In the Kondo regime or after the double bound mode decays i.e. $t\sim\gamma^{-1}$, the noise slows the leakage of magnetization and leads to a quantum Zeno effect as the dephasing rate increases. In the long-time limit, independent of the boundary coupling $J$, the impurity magnetization falls off as $S^{z}_{0}(t)\sim e^{-t/\tau}$ with $\tau\propto\frac{1}{\gamma L^{2}}$ in both regimes. This behavior of the impurity in the Kondo regime is consistent with the heating of the bulk induced by Hermitian Lindblad operators [94, 95, 96], driving the impurity to (asymptotically free) weak coupling behavior [97, 98, 99, 100].
@@
Figure 4: Data collapse for $S^{z}(x,t)=S^{z}(\xi)$ with $\xi=xt^{-\frac{1}{2}}$ in the XX case (filled markers, with system size $L=40$) and XXX case (hollow markers, with system size $L=10$) starting from a domain wall initial state (as shown in the color plot in the right inset for XX case with $\gamma=2$) under various dephasing rates $\gamma$. Different marker shapes correspond to the spin configuration at different time slices, and different marker color refers to different dephasing rates. The left inset shows the dependence of the diffusion constant $D$ on $\gamma$ (filled markers for XX case and hollow for XXX case), and the dashed lines are the analytical result for the noisy XX chain $D=\frac{1}{2\gamma}$.